A clean glass capillary tube, of internal diameter 0.04 cm, is held vertically with its lower end below the surface of clean water in a beaker, and with 10 cm of the tube above the surface. What will happen if the tube is depressed until only 5 cm of its length is above the surface? The surface tension of water is 7.2 x 10-2 Nm-1.
1
Expert's answer
2020-10-29T07:00:45-0400

Explanations & Calculations


  • The relationship between the height a particular liquid rises in a given capillary tube due to the surface tension is given by

"\qquad\qquad
\begin{aligned}
\small h &= \small \frac{2T\cos\theta }{r\rho g}\\

\end{aligned}" : "\small \theta" is the contact angle between water & glass


  • And for the given situation (10cm above the water level) the capillary rise is

"\qquad\qquad
\begin{aligned}
\small h_1 &= \small \frac{2\times 0.072Nm^{-1}}{0.0002m \times 10^3kgm^{-3}\times 9.8ms^{-2}} \cos\theta\\
\small &= \small 0.0735\cos\theta \,\,m \\
\small&=\small 7.35\cos \theta \,\,cm
\end{aligned}"

  • It's obvious that the capillary rise will not change as none of the components in the equation is changed even after the tube is depressed further.


  • After it is depressed only 5 cm is above the water & if "\small h_1<5cm" no water will spill & if "\small h_1>>5cm" water will spill.
  • If the capillary rise just a very little more that 5cm water will stay in the tube until the contact angle reaches 180 degrees without spill.